OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. m = 2, A(_, 0), B(5, 10)

OpenStudy (imstuck):

Slope formula is this:\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]So fill in what you know using the slope as "m" in your formula. You are given x2, y1, and y2, and need to find x1. So do it like this:

OpenStudy (imstuck):

\[2=\frac{ 10-0 }{ 0-x }\]10 divided by what number will give you a positive 2 as a result? If you put a 5 in there you will get 10/ -5 = -2, so the number you need to fill in for x has to be a negative number because a negative minus a negative gives you a positive. So it is -5. Here's proof:

OpenStudy (imstuck):

\[2=\frac{ 10-0 }{ 0-(-5) }->2=\frac{ 10-0 }{ 0+5 }->2=\frac{ 10 }{ 5 }\]See?

OpenStudy (imstuck):

So your missing blank is -5.

OpenStudy (anonymous):

so it would be 5 ok -5